OpenStudy (anonymous):

determine the average rate of change of f(x)=x^4-2x between the given values x=-1,x=3

OpenStudy (anonymous):

Do you know it's formula ?

OpenStudy (anonymous):

yeah i was tryin to work it out but i dont know what to do with the two x values

OpenStudy (anonymous):

\[\Large Average rate of change = \frac{f(x+h)-f(x)}{h}\] where x=x1 , h= x2-x1 Got it ?

OpenStudy (anonymous):

ok let me try it

OpenStudy (anonymous):

i got 29 but its telling me its the wrong answer

OpenStudy (anonymous):

so for x would be -1 and h would be 4 right

OpenStudy (anonymous):

F(x)=x^4-2x . (x1=-1) ,(x2=-2) Substitute : \[\Large \frac{(x^4-2x)-(x^4-2x)}{x_{2}-x_{1}}\] \[\Large \frac{((-2)^4-2(-2))-((-1)^4-2(-1))}{-2-(-1)}\] The result will be : -17 . Sry ,There is lag :s

OpenStudy (anonymous):

how did you get negative 2

OpenStudy (anonymous):

its the value of x2

OpenStudy (anonymous):

i dont know why but it keeps saying thats not the answer ill see what i can do thanks either way

OpenStudy (anonymous):

do you think i have to take the derivative first

OpenStudy (anonymous):

The rate of change is the first degree of differential

OpenStudy (anonymous):

OMG Dang ,Sry i thought that the value of x2 = -2 :OOOOOOOOOOOOOO that is soo weird I'am sry Look we will do the same steps but instead of -2 we will put 3 .

OpenStudy (anonymous):

okay let me try it

OpenStudy (anonymous):

yeah that was it thank you so much :)

OpenStudy (anonymous):

Yw,and sry for that i was kinda hyper :P